# Fact-Aware Multimodal Retrieval Augmentation for Accurate Medical Radiology Report Generation

### v1

@@ Abstract
Multimodal foundation models hold significant potential for automating radiology report generation, thereby assisting clinicians in diagnosing cardiac diseases. However, generated reports often suffer from serious factual inaccuracy. In this paper, we introduce a fact-aware multimodal retrieval-augmented pipeline in generating accurate radiology reports (FactMM-RAG). We first leverage RadGraph to mine factual report pairs, then integrate factual knowledge to train a universal multimodal retriever. Given a radiology image, our retriever can identify high-quality reference reports to augment multimodal foundation models, thus enhancing the factual completeness and correctness of report generation. Experiments on two benchmark datasets show that our multimodal retriever outperforms state-of-the-art retrievers on both language generation and radiology-specific metrics, up to 6.5% and 2% score in F1CheXbert and F1RadGraph. Further analysis indicates that employing our factually-informed training strategy imposes an effective supervision signal, without relying on explicit diagnostic label guidance, and successfully propagates fact-aware capabilities from the multimodal retriever to the multimodal foundation model in radiology report generation. [^1] Code will be public upon acceptance.**footnotetext: These authors contributed equally to this work.
@@ 2 Related Work
One closely related line of work to ours is retrieval-based radiology report generation given only radiology images. For instance, Li et al. (2018) proposes a retrieval policy module to update radiology reports via hierarchical reinforcement learning; Endo et al. (2021) employs image-text embeddings from contrastive learning for retrieval-augmented radiology report generation; Ramesh et al. (2022) proposes synthesizing additional reports and reducing hallucinations from reference report priors to improve report generation.
@@ 3.1 Fact-aware Multimodal Retrieval
$s(q_{txt},d_{txt})=\frac{2\cdot(\hat{q}_{txt}\cap\hat{d}_{txt})}{\text{length}(\hat{q}_{txt})\cdot\text{length}(\hat{d}_{txt})},$ (1)
where $\hat{q}_{txt},\hat{h}_{txt}$ denotes reports with only annotated entities and relations in RadGraph structured form. We then set a strict threshold $\delta$ to filter out searched reports with low similarity score:
@@ 4 Experimental Setup
Evaluation Metrics. We evaluate our proposed system using both natural language generation and medically-tailored evaluation metrics.
For language fluency measures, we use ROUGE-L Lin (2004) to evaluate the longest common subsequence overlap between the generated and reference findings, and BERTScore Zhang et al. (2020a) to evaluate non-clinical semantic sentence similarity.
For clinical accuracy measures, we employ CheXbert Smit et al. (2020) to generate the ground-truth diagnostic labels for finding reports, identifying 14 different types of observations. Following Delbrouck et al. (2023), we then calculate the F1CheXbert (Zhang et al., 2020b), which is the F1-score for 5 observations (Cardiomegaly, Edema, Consolidation, Atelectasis, Pleural Effusion) by comparing the generated report with the reference report’s classifications. Beyond using the limited diagnostic labels for evaluation, we also adopt F1RadGraph (Jain et al., 2021) to measure factual correctness by calculating the overlap in radiological entities and clinical relations between the generated report and the reference report.
@@
Implementation Details. In our experiments, we use MARVEL (Zhou et al., 2024) as our multimodal retriever backbone. MARVEL is a language model based on T5-ANCE (Yu et al., 2023a), trained with modality-balanced hard negatives. We use LLaVA Liu et al. (2023) as our multimodal foundation model backbone. Since each radiology study contains multiple image views for each patient, we select the frontal view. We also concatenate the finding and impression sections to form the X-ray report. To reduce training costs and address factual report pair imbalances, we rerank the retrieved reports by factual similarity and use the top 2 factual report pairs for each query to train our multimodal retriever. We leave the training details in Appendix A.
@@
(d) F1CheXbert Threshold: 1.0
Figure 2: Factual performance of FactMM-RAG controlled by different F1CheXbert and F1RadGraph thresholds.
@@ 5 Evaluation Results
In this section, we present our experimental results. We first evaluate the overall performance between different retrievers under two settings in section 5.1. Next, we discuss the ablation studies in section 5.2. We then explore the fact-aware capability of our retriever in section 3 and section 5.4. Lastly, we show the superiority of our retriever through a case study in section 5.5.
@@ 5.3 Fact-aware Capability Control
Figure 3: Retrieval evaluation of FactMM-RAG with different F1CheXbert and F1RadGraph thresholds.
@@
Table 3: One case study from MIMIC-CXR. Cyan text indicates radiological consistency with the ground-truth report. Orange text highlights extra accurate details provided by FactMM-RAG compared to Med-MARVEL. Red text denotes observations missing in Med-MARVEL.
@@ 5.4 Fact-aware Capability Propagation
Figure 4: Analysis of fact-aware capability propagation evaluated by MRR retrieval metric.
@@ 5.5 Case Study
In this section, we present two examples from MIMIC-CXR to qualitatively analyze our retriever’s fact-aware capability, as illustrated in Table 3. In the first example, we observe that FactMM-RAG provides symptom observations consistent with the ground-truth report and generates more accurate factual details compared to Med-MARVEL, e.g., “post median sternotomy, atelectasis, not pneumothorax”; In the second example, we further observe that although both retrievers generate reports with diagnostic labels matching the ground-truth report, FactMM-RAG provides additional details compared to Med-MARVEL, such as “pulmonary vasculature is normal, no acute osseous abnormalities”. These characteristics confirm that adopting our fact-aware retriever can assist multimodal foundation models in generating more accurate radiology reports.
@@ 6 Conclusion
In this paper, we aim at improving radiology report generation by introducing a fact-informed medical multimodal retriever for retrieval-augmented generation. In particular, we utilize RadGraph to annotate chest radiograph reports and mine clinically-relevant pairs. We integrate factual information into a universal multimodal retriever, presenting FactMM-RAG, a fact-aware multimodal retrieval-augmented radiology report generation pipeline. FactMM-RAG outperforms all state-of-the-art retrievers evaluated by factual correctness and textual coherence for final report generation in MIMIC-CXR and CheXpert datasets. We further confirm the benefit of our multimodal retriever from the analysis of fact-aware capability control and propagation. Given the pervasive applications of machine learning in clinical diagnoses using chest X-rays, we hope our factual-informed approach inspires further work in multimodal generative artificial intelligence in healthcare contexts.
@@ 7 Limitations
Another concern lies in the chosen evaluation metrics, F1RadGraph and F1CheXbert. F1CheXbert reflects high-level observational accuracy, while F1RadGraph assesses the correctness of radiology entities and clinical relationships. However, other radiologically-specific metrics, such as report conciseness and clarity, should also be considered (Sureka et al., 2014). Ideally, we should incorporate methods of evaluation directly aligned with human evaluations or involve domain expertise itself in our pair-mining and final evaluation procedure.

### v2

@@ Abstract
Multimodal foundation models hold significant potential for automating radiology report generation, thereby assisting clinicians in diagnosing cardiac diseases. However, generated reports often suffer from serious factual inaccuracy. In this paper, we introduce a fact-aware multimodal retrieval-augmented pipeline in generating accurate radiology reports (FactMM-RAG). We first leverage RadGraph to mine factual report pairs, then integrate factual knowledge to train a universal multimodal retriever. Given a radiology image, our retriever can identify high-quality reference reports to augment multimodal foundation models, thus enhancing the factual completeness and correctness of report generation. Experiments on two benchmark datasets show that our multimodal retriever outperforms state-of-the-art retrievers on both language generation and radiology-specific metrics, up to 6.5% and 2% score in F1CheXbert and F1RadGraph. Further analysis indicates that employing our factually-informed training strategy imposes an effective supervision signal, without relying on explicit diagnostic label guidance, and successfully propagates fact-aware capabilities from the multimodal retriever to the multimodal foundation model in radiology report generation. [^1] Our code is available at https://github.com/cxcscmu/FactMM-RAG**footnotetext: These authors contributed equally to this work.
@@ 2 Related Work
One closely related line of work to ours is retrieval-based radiology report generation given only radiology images. For instance, Li et al. (2018) proposes a retrieval policy module to update radiology reports via hierarchical reinforcement learning; Endo et al. (2021) employs image-text embeddings from contrastive learning for retrieval-augmented radiology report generation; Ramesh et al. (2022) proposes synthesizing additional reports and reducing hallucinations from reference report priors to improve medical radiological report generation.
@@ 3.1 Fact-aware Multimodal Retrieval
$s(q_{txt},d_{txt})=\frac{2\cdot(\hat{q}_{txt}\cap\hat{d}_{txt})}{\text{length}(\hat{q}_{txt})+\text{length}(\hat{d}_{txt})},$ (1)
where $\hat{q}_{txt},\hat{d}_{txt}$ denotes reports with only annotated entities and relations in RadGraph structured form. We then set a strict threshold $\delta$ to filter out searched reports with low similarity score:
@@ 4 Experimental Setup
Evaluation Metrics. We evaluate our proposed system using both natural language generation and medically-tailored evaluation metrics. For language fluency measures, we use ROUGE-L Lin (2004) to evaluate the longest common subsequence overlap between the generated and reference findings, and BERTScore Zhang et al. (2020a) to evaluate non-clinical semantic sentence similarity.
For clinical accuracy measures, we use CheXbert Smit et al. (2020) to generate the ground-truth diagnostic labels for finding reports. Following Delbrouck et al. (2023), we then calculate the F1CheXbert (Zhang et al., 2020b), which is the F1-score for 5 observations (Cardiomegaly, Edema, Consolidation, Atelectasis, Pleural Effusion) by comparing the generated report with the reference report’s classifications. Beyond using the limited diagnostic labels for evaluation, we also adopt F1RadGraph (Jain et al., 2021) to measure factual correctness by calculating the overlap in radiological entities and clinical relations between the generated report and the reference report. See Appendix A.3 for more details.
@@
We also compare our method with non-RAG approach. "No Retriever" refers to directly fine-tuning the backbone to generate reports without retrieval augmentation; ORGan (Hou et al., 2023) first creates an observation plan, then feeds the plan and radiographs to generate the report through tree reasoning mechanism. Upper-bound results using an oracle in training corpus with top-1 factual similarity to test query report are presented. More details are in Appendix A.3.
Implementation Details. In our experiments, we use MARVEL (Zhou et al., 2024) as our multimodal retriever backbone. MARVEL is a language model based on T5-ANCE (Yu et al., 2023a), trained with modality-balanced hard negatives. We use LLaVA Liu et al. (2023) as our multimodal foundation model backbone. Since each radiology study contains multiple image views for each patient, we select the frontal view. We also concatenate the finding and impression sections to form the X-ray report. To reduce training costs and address factual report pair imbalances, we rerank the retrieved reports by factual similarity and use the top 2 factual report pairs for each query to train our multimodal retriever. We leave more training details in Appendix A.1 and A.2.
@@
(d) F1CheXbert Threshold: 1
Figure 2: Factual performance of FactMM-RAG controlled by different F1CheXbert and F1RadGraph thresholds. We vary the F1RadGraph thresholds under one fixed F1CheXbert threshold selected from {0, 0.4, 0.8, 1}.
@@ 5 Evaluation Results
In this section, we present our experimental results. We first evaluate the overall performance between different retrievers in section 5.1. Next, we discuss the ablation studies in section 5.2. We then explore the fact-aware capability of our retriever in section 5.3 and section 5.4. Lastly, we show the superiority of our retriever through a case study in section 5.5.
@@ 5.1 Overall Performance
To establish the effectiveness of our RAG approach, we also show that FactMM-RAG significantly outperforms the fine-tuned backbone without retrieval augmentation by 10% and achieves competitive improvement over the strong non-RAG ORGan baseline.
@@ 5.3 Fact-aware Capability Control
Figure 3: Retrieval evaluation of FactMM-RAG with different F1CheXbert and F1RadGraph thresholds. MRR calculates the mean reciprocal of rank at which the first relevant report that meets two factual similarity thresholds with query report is retrieved.
@@
Table 3: Case study on generated reports from MIMIC-CXR. Cyan text indicates radiological consistency with the ground-truth report. Orange text highlights extra accurate details provided by FactMM-RAG compared to Med-MARVEL. Red text denotes observations missing in Med-MARVEL.
@@ 5.4 Fact-aware Capability Propagation
Figure 4: Analysis of fact-aware capability propagation. The $x$-axis MRR measures the retriever’s performance on retrieving factually relevant reports.
@@ 5.5 Case Study
In this section, we present two examples from MIMIC-CXR to qualitatively analyze our retriever’s fact-aware capability, as illustrated in Table 3. In the first example, we observe that FactMM-RAG provides symptom observations consistent with the ground-truth report and generates more accurate factual details compared to Med-MARVEL, e.g., “post median sternotomy, atelectasis, not pneumothorax”; In the second example, we further observe that although both retrievers generate reports with diagnostic labels matching the ground-truth report, FactMM-RAG provides additional details compared to Med-MARVEL, such as “pulmonary vasculature is normal, no acute osseous abnormalities”. These characteristics confirm that adopting our fact-aware retriever can assist multimodal foundation models in generating more accurate radiology reports. We show retrieved reports from two samples in Appendix A.4.
@@ 6 Conclusion
In this paper, we aim at improving radiology report generation by introducing a fact-informed medical multimodal retriever for retrieval-augmented generation. In particular, we utilize RadGraph to annotate chest radiograph reports and mine clinically-relevant pairs. We integrate factual information into a universal multimodal retriever, presenting FactMM-RAG, a fact-aware multimodal retrieval-augmented radiology report generation pipeline. FactMM-RAG outperforms all state-of-the-art retrievers evaluated by factual correctness and textual coherence for final report generation in MIMIC-CXR and CheXpert datasets. We further confirm the benefit of our multimodal retriever from the analysis of its fact-aware capability.
@@ 7 Limitations
Another concern lies in the chosen evaluation metrics, F1RadGraph and F1CheXbert. F1CheXbert reflects high-level observational accuracy, while F1RadGraph assesses the correctness of radiology entities and clinical relationships. However, other radiologically-specific metrics, such as report conciseness and clarity, should also be considered (Sureka et al., 2014). Ideally, we should incorporate methods of evaluation directly aligned with human evaluations or involve domain expertise itself in our pair-mining and final evaluation procedure. Moreover, it is worth performing a long-tail evaluation by leveraging more fine-grained ground-truth label annotations (Holste et al., 2023).
8 Ethics Considerations
A key ethical consideration in our work is the use of de-identified, credentialed medical data. In particular, the responsible usage policy of the MIMIC-CXR dataset strictly prohibits sharing access with third parties. To gain access to MIMIC-CXR, we completed a training course and signed a data use agreement, ensuring compliance with patient privacy regulations.
@@ A.2 RAG Finetuning Procedure
Radiology Image Med-Marvel FactMM-RAG Reference
[Uncaptioned image] A single portable chest radiograph was obtained. Bilateral pleural effusions and mild atelectasis have increased. Cardiomegaly is unchanged. There is no consolidation or pneumothorax. Pacing leads, sternotomy wires, vascular clips, and abdominal surgical clips are unchanged. No focal consolidation is identified. There is unchanged appearance of opacifications in the left lung base, likely due to a combination of atelectasis and pleural effusion. There is a small right pleural effusion. Mild pulmonary edema persists. The heart is moderately enlarged, but stable. Left sided pacemaker is seen with transvenous leads in the right atrium, right ventricle, and left ventricle. The patient is status post median sternotomy and CABG. Left-sided dual-chamber pacemaker is noted with leads terminating in right atrium and right ventricle, unchanged. Cardiomegaly is similar. There is continued mild to moderate pulmonary edema, slightly improved compared to the prior exam. Small layering bilateral pleural effusions also may be slightly decreased in the interval. Bibasilar airspace opacities likely reflect atelectasis.There is no pneumothorax. No acute osseous abnormalities are visualized.
F1RadGraph 0.274 0.345
CheXbert Observations Cardiomegaly, Atelectasis, Pleural Effusion Cardiomegaly, Edema, Atelectasis, Pleural Effusion Cardiomegaly, Edema, Atelectasis, Pleural Effusion
[Uncaptioned image] The heart is mildly enlarged with a left ventricular configuration. There is mild-to-moderate unfolding of the thoracic aorta. The arch is partly calcified. The mediastinal and hilar contours appear unchanged. There are streaky left basilar opacities suggesting minor atelectasis. A small eventration is noted along the anterior right hemidiaphragm. There is an air-fluid level in the stomach. Air-fluid levels are seen in the epigastric region. There is no evidence for free air. Cholecystectomy clips project over the right upper quadrant. Moderate degenerative changes are similar along the mid thoracic spine. Moderate enlargement of the cardiac silhouette is unchanged. The aorta remains tortuous. The mediastinal and hilar contours are normal. Pulmonary vasculature is normal. Streaky atelectasis is noted in the left lower lobe. The right lung is clear. No focal consolidation, pleural effusion or pneumothorax is present. Multiple clips are noted within the left upper abdomen. Moderate enlargement of the cardiac silhouette with a left ventricular predominance is unchanged. The aorta remains tortuous, and the hilar contours are stable. Pulmonary vascularity is not engorged. There is minimal atelectasis within the lung bases, but no focal consolidation is present. No pleural effusion or pneumothorax is identified. There are no acute osseous abnormalities.
F1RadGraph 0.197 0.621
Table 4: Case study on retrieved reports from MIMIC-CXR. Cyan text indicates radiological consistency with the ground-truth report. Orange text highlights extra accurate details provided by FactMM-RAG compared to Med-MARVEL. Red text denotes observations missing in Med-MARVEL.
A.3 Evaluation Details
Here, we provide implementation details regarding the evaluation methodology.
F1-RadGraph. For F1-RadGraph score computation, we follow previous work (MIMIC-CXR-RRS)  [^2] https://vilmedic.app/papers/acl2023/ in employing $\textrm{RG}_{ER}$ as F1-Radgraph score computation on an instance level. Using the radgraph library implementation, this equates to utilizing reward_level="partial".
F1-CheXbert. F1-CheXbert score computation consists of the micro-averaged F1-score between 5 selected classes from the CheXbert labeler. Naturally, F1-CheXbert scores are only computable over entire datasets. For instance-level CheXbert scores (used for pair mining), we employ the proportion of equivalent predicted classes between a reference and predicted text sample. These instance-level F1-CheXbert scores can be computed using np.sum(ref == hyp) / 5, and take on values $\in\{0.0,0.2,0.4,0.6,0.8,1.0\}$.
CheXpert Hidden Test Set. We use the 1000 hidden test reports from MIMIC-CXR-RRS and download the CheXpert images from Stanford AIMI Shared Datasets  [^3] https://stanfordaimi.azurewebsites.net/datasets/8cbd9ed4-2eb9-4565-affc-111cf4f7ebe2.
Oracle Retrieval. Oracle Retrieval is performed via ground-truth access to a reference document’s generated report. For training queries, this is always known, and an oracle retriever would obtain documents as $Oracle(q_{i})\doteq\underset{j\in\textrm{corpus},j\neq i}{\arg\max}~{}s(q_{i},d_{j})$, where $s(q,d)$ is the sum of the F1-RadGraph and F1-CheXbert instance-wise scores. In practice, this results in retrieving samples with F1-CheXbert=$1.0$ and the largest F1-RadGraph score within the partition. Test-time retrieval performs the same operation, without the restriction of $j\neq i$ as self-retrieval is not possible due to the corpus being the training dataset.
Oracle RAG. Oracle-LLaVA is obtained by fine-tuning LLaVA under identical conditions, utilizing Oracle Retrieval for retrieving documents in the training and test set.
A.4 Case study on Retrieved Reports
We now conduct case study on the retrieved reports shown in Table . We show that our FactMM-RAG captures most of the factual details in retrieved reports compared to the ground-truth reports. Thus, the factual correctness of our retriever can be propagated to the multimodal foundation models effectively. However, the reports retrieved from Med-MARVEL contain erroneous information, which negatively impacts the report generation by multimodal foundation models.